OpenStudy (anonymous):

If a resistance of R ohms is connected across a battery of E volts with internal resistance (the resistance within the battery) r ohms, then the power in watts dissipated in the external resistor is given by P = (E^2R)/((R+r)^2. If E and r are fixed, find R so that the power dissipated externally is a maximum.

OpenStudy (wmj259):

Physics 2?

OpenStudy (anonymous):

Calculus 1

OpenStudy (wmj259):

Interesting*

OpenStudy (anonymous):

I found an answer but I do not know if it is right.

OpenStudy (anonymous):

\[R = \frac{ 1 }{ 2 }E ^{2} - \frac{ 1 }{ 2 }r - \frac{ 1 }{ 2 }\sqrt{E ^{4}-6rE ^{2}+r ^{2}}\]

OpenStudy (anonymous):

@jim_thompson5910 @jdoe0001

OpenStudy (anonymous):

@Zarkon @zepdrix

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@geerky42